OpenStudy (anonymous):

A 1.5-k Ω resistor is in series with unknown resistance. the applied voltage, equals 36 V and the series current is 14.4 mA. Calculate the value of the unknown resistor.

OpenStudy (anonymous):

Solve for k: -k+1.5 A = 0 Factor the left hand side. Factor constant terms from the left hand side: 1.5 (-0.666667 k+1. A) = 0 Divide both sides by a constant to simplify the equation. Divide both sides by 1.5: -0.666667 k+1. A = 0 Isolate terms with k to the left hand side. Subtract 1. A from both sides: -0.666667 k = -1. A Solve for k. Divide both sides by -0.666667: Answer: | | k = 1.5 A